Find all two-digit numbers that give a sum of 154 when the original number and the number obtained by reversing its digits are added.
Answer by stanbon(75887)

Find all two-digit numbers that give a sum of 154 when the original number and the number obtained by reversing its digits are added.
-----
Let the original number be 10t+u
--
Reverse number is 10u+t
-----
Equation:
10t+u + 10u+t = 154
11t + 11u = 154
t + u = 14
----
Both t and u must be single digit numbers
Pairs that work are 5/9,6/8,7/7
---
Original numbers are 59 or 68 or 77
------------------------------
Cheers,
Stan H.
==========